# Accuracy is Not Agreement: Expert-Aligned Evaluation of Crash Narrative Classification Models

### v1

@@ Abstract
This study investigates the relationship between deep learning (DL) model accuracy and expert agreement in classifying crash narratives. We evaluate five DL models—including BERT variants, USE, and a zero-shot classifier—against expert labels for 300 narratives, and extend the analysis to four large language models (LLMs): GPT-4, LLaMA 3, Qwen, and Claude. Our findings reveal an inverse relationship: models with higher technical accuracy often show lower agreement with human experts, while LLMs demonstrate stronger expert alignment despite lower accuracy. We use Cohen’s Kappa and Principal Component Analysis (PCA) to quantify and visualize model-expert agreement, and employ SHAP analysis to explain misclassifications. Results show that expert-aligned models rely more on contextual and temporal cues, rather than location-specific keywords. These findings suggest that accuracy alone is insufficient for safety-critical NLP tasks. We argue for incorporating expert agreement into model evaluation frameworks and highlight the potential of LLMs as interpretable, cost-effective tools in crash analysis pipelines.
@@ I Introduction
Road traffic crashes remain a major global public health issue, with U.S. fatal crashes rising 9.9% in 2021 despite vehicle safety improvements [1]. To better understand crash dynamics, researchers increasingly rely on crash databases that combine structured data with unstructured crash narratives—free-text descriptions from responding officers that offer contextual insight often missing from coded fields. These narratives have been effectively used to study various crash types, including work zones [2], military collisions [3], and distraction-related incidents [4]. Prior work also shows that narrative-derived features can uncover hidden risk factors [5] and support scenario reconstruction [6]. Manual review of narratives, however, is labor-intensive and subjective, limiting scalability. To address this, researchers have increasingly adopted natural language processing (NLP) and deep learning (DL) for classification tasks [7]. While video-based multimodal approaches show promise [8, 9], text-based models remain more feasible due to lower resource demands and broad availability of narrative data [10]. Despite their technical promise, DL models often lack interpretability and can produce outputs that diverge from expert judgment, especially problematic in safety-critical contexts. Studies in work zones and bicycle-related crashes [2, 11, 12] show that high accuracy alone does not ensure practical utility. Moreover, narrative accuracy is limited by reporting inconsistencies from officers or data entry errors [7, 13]. This gap between model performance and expert interpretation risks real-world deployment. In this work, we propose an evaluation framework that complements accuracy with expert agreement metrics such as Cohen’s Kappa and applies Principal Component Analysis (PCA) to visualize model-expert divergence, offering a more robust validation pathway for automated crash classification systems.
@@ II-A Data Collection and Preprocessing
This study utilizes crash report narratives from the Iowa Department of Transportation for 2019-2020, totaling 55,081 records with 37,565 having non-empty narrative fields. The narratives describe crash events in natural language format, written by law enforcement officials. The average narrative contains 38 words, with a high standard deviation (25.7) indicating significant length variation. The ground-truth labels (intersection-related vs non-intersection-related) were derived from the crash report’s coded fields, particularly the ”location” attribute.
@@
The crash narrative dataset contained a ’crash key’ column and five narrative fields (Narrative 1 through Narrative 5), which we combined to generate comprehensive narratives for each crash key. The complete dataset included 100,812 crash data records for 2019 and 72,700 for 2020, 57,479 crash narrative records for 2019 and 36,888 for 2020. The dataset comprised 173,512 crash data records and 94,367 crash narrative records across both years (2019-2020).
@@ 3 Non-intersection: Railroad grade crossing
4-7, 96 Other Non-intersection types (Crossover-related)
@@ II-B Expert Annotations
The annotation process followed a rigorous protocol to minimize bias, including standardized guidelines, randomized narrative presentation, independent classification, and confidence ratings for ambiguous cases. We analyzed labeling patterns across different affiliations to assess potential biases, finding no statistically significant differences ($\chi^{2}=2.14$, $p=0.14$).
@@
The inter-expert agreement demonstrated moderate consistency ($\kappa=0.58$, 95% CI: 0.49–0.65), underscoring the inherent subjectivity in interpreting crash narratives. This level of agreement provides a critical benchmark, as it is unlikely that any automated system would outperform the consensus reached by human experts.
@@ IV-B Expert Agreement Analysis
Inter-rater agreement helps determine whether experts’ opinions align when assessing data. As shown in Table IV, our analysis revealed a surprising inverse relationship between model accuracy and expert agreement. Among specialized DL models, USE and BZS showed the highest agreement percentage with experts (89.80%), while BSE showed the least (81.63%) despite having the highest technical accuracy (83.00%). When quantified using Cohen’s Kappa, BSE demonstrated the weakest agreement with experts ($\kappa=0.41$, 95% CI: 0.35–0.47), whereas USE ($\kappa=0.67$, 95% CI: 0.58–0.74) and BZS ($\kappa=0.63$, 95% CI: 0.55–0.71) showed substantially higher alignment with expert classifications.
@@ IV-E Model-Expert Discrepancy Analysis
Our SHAP analysis (Figure 5) revealed that models often placed excessive weight on specific location-related terms without adequately considering the complete contextual narrative. We systematically quantified feature importance across all models, as shown in Table VI. The BSE model showed the highest reliance on direct intersection-related keywords, with terms like ”intersection” and ”signal” accounting for 42.6% of its total feature importance. In contrast, USE distributed feature importance more evenly, with the top five keywords accounting for only 27.3% of its total attribution, placing substantial weight on contextual modifiers like ”approaching” (7.2%), ”after” (6.1%), and ”before” (5.8%).
@@ IV-F1 LLM Evaluation Methodology
Each LLM was provided with identical prompts asking them to classify crash narratives as intersection-related or non-intersection-related. The prompt included general instructions about crash location classification, a brief explanation of intersection criteria, and the narrative text. To ensure fairness, we used the following standardized prompt structure:
@@ V Discussion and Conclusion
This study demonstrates an inverse relationship between deep learning model accuracy and expert judgment alignment in crash narrative classification. Our analysis of five specialized DL models and four LLMs reveals that models with higher technical accuracy often show lower agreement with expert classifications, while models with moderate accuracy better capture human reasoning patterns.
USE and BZS demonstrated more substantial expert alignment among specialized models, suggesting sentence-level embeddings better reflect holistic expert reasoning than BERT-based approaches that excel at localized pattern recognition but miss contextual nuance. LLMs achieved substantially higher expert agreement ($\kappa$ up to 0.81) despite lower technical accuracy, with their high inter-model consensus (94.3%) indicating convergence on interpretations prioritizing contextual understanding over keyword recognition.
The observed discrepancies stem from fundamental differences in decision-making: experts employ holistic reasoning considering temporal sequences and implicit spatial relationships, while high-accuracy models rely heavily on location-specific terms. SHAP analysis confirmed that models with higher expert agreement distribute feature importance more evenly across contextual terms rather than fixating on keywords.
For transportation agencies, these findings suggest practical implementation strategies: (1) incorporate expert-alignment metrics alongside accuracy in model evaluations, (2) implement hybrid approaches leveraging complementary model strengths, (3) establish confidence thresholds to route uncertain predictions for human review, and (4) consider LLMs as cost-effective intermediaries for model validation.
Future research should explore whether this inverse accuracy-agreement relationship persists across different classification tasks, develop approaches combining high-accuracy models with expert-aligned LLMs, and investigate methods to distill contextual reasoning capabilities into computationally efficient systems. As automated systems increasingly support transportation safety decisions, ensuring alignment between algorithmic classifications and expert judgment remains essential.

### v2

@@ Abstract
This study investigates the relationship between deep learning (DL) model accuracy and expert agreement in classifying crash narratives. We evaluate five DL models—including BERT variants, USE, and a zero-shot classifier—against expert labels and narratives, and extend the analysis to four large language models (LLMs): GPT-4, LLaMA 3, Qwen, and Claude. Our findings reveal an inverse relationship: models with higher technical accuracy often show lower agreement with human experts, while LLMs demonstrate stronger expert alignment despite lower accuracy. We use Cohen’s Kappa and Principal Component Analysis (PCA) to quantify and visualize model-expert agreement, and employ SHAP analysis to explain misclassifications. Results show that expert-aligned models rely more on contextual and temporal cues than location-specific keywords. These findings suggest that accuracy alone is insufficient for safety-critical NLP tasks. We argue for incorporating expert agreement into model evaluation frameworks and highlight the potential of LLMs as interpretable tools in crash analysis pipelines.
@@ I Introduction
Road traffic crashes remain a major global public health issue, with U.S. fatal crashes rising 9.9% in 2021 despite vehicle safety improvements [1]. To better understand crash dynamics, researchers increasingly rely on crash databases that combine structured data with unstructured crash narratives—free-text descriptions from responding officers that offer contextual insight often missing from coded fields. These narratives have been effectively used to study various crash types, including work zones [2], military collisions [3], and distraction-related incidents [4]. Prior work also shows that narrative-derived features can uncover hidden risk factors [5] and support scenario reconstruction [6]. Manual review of narratives, however, is labor-intensive and subjective, limiting scalability. To address this, researchers have increasingly adopted natural language processing (NLP) and deep learning (DL) for classification tasks [7]. While video-based multimodal approaches show promise [8, 9], text-based models remain more feasible due to lower resource demands and broad availability of narrative data [10]. Despite their technical promise, DL models often lack interpretability and can produce outputs that diverge from expert judgment, which is especially problematic in safety-critical contexts. Studies in work zones and bicycle-related crashes [2, 11, 12] show that high accuracy alone does not ensure practical utility. Moreover, narrative accuracy is limited by reporting inconsistencies from officers or data entry errors [7, 13]. This gap between model performance and expert interpretation poses a risk to real-world deployment. In this work, we propose an evaluation framework that complements accuracy with expert agreement metrics such as Cohen’s Kappa and applies Principal Component Analysis (PCA) to visualize model-expert divergence, offering a more robust validation pathway for automated crash classification systems.
@@ II-A Data Collection and Preprocessing
This study utilizes the Iowa Department of Transportation crash report narratives for 2019-2020. The narratives describe crash events in natural language format, written by law enforcement officials. The average narrative contains 38 words, with a high standard deviation (25.7) indicating significant length variation. The ground-truth labels (intersection-related vs non-intersection-related) were derived from the crash report’s coded fields, particularly the ”location” attribute. The crash narrative dataset contained a ’crash key’ column and five narrative fields (Narrative 1 through Narrative 5), which we combined to generate comprehensive narratives for each crash key. The complete dataset included 100,812 crash data records for 2019 and 72,700 for 2020, 57,479 crash narrative records for 2019 and 36,888 for 2020. The dataset comprised 173,512 crash data records and 94,367 crash narrative records across both years (2019-2020).
@@ 3 Non-intersection: Railroad grade crossing
4-7 Other Non-intersection types (Crossover-related)
@@ II-B Expert Annotations
The annotation process followed a rigorous protocol to minimize bias, including standardized guidelines, randomized narrative presentation, independent classification, and confidence ratings for ambiguous cases. We analyzed labeling patterns across different affiliations to assess potential biases and found no statistically significant differences ($\chi^{2}=2.14$, $p=0.14$).
@@
The inter-expert agreement demonstrated fair consistency ($\kappa=0.25$, 95% CI: 0.18–0.32), with pairwise agreement averaging approximately 62.5%. The Fleiss’ Kappa value of 0.20 (95% CI: 0.14–0.26) indicates slight agreement across all four experts, reflecting the challenge of achieving consensus in crash narrative interpretation. These values align with the ground-truth labels agreement percentages (48.98% to 67.35%, see Table IV), suggesting that potential noise in structured data fields contributes to variability. This level of agreement provides a benchmark for evaluating automated systems against expert judgment.
@@ IV-B Expert Agreement Analysis
As shown in Table IV, our analysis revealed a surprising inverse relationship between model accuracy and expert agreement. Among specialized DL models, USE and BZS showed the highest agreement with experts (89.80%), while BSE showed the least (81.63%) despite having the highest technical accuracy (83.00%). When quantified using Cohen’s Kappa, BSE demonstrated the weakest agreement with experts ($\kappa=0.41$, 95% CI: 0.35–0.47), whereas USE ($\kappa=0.67$, 95% CI: 0.58–0.74) and BZS ($\kappa=0.63$, 95% CI: 0.55–0.71) showed substantially higher alignment with expert classifications. Human experts achieved ground-truth labels agreement ranging from 48.98% to 67.35%, consistent with the fair inter-expert Kappa of 0.25 (95% CI: 0.18–0.32) reported in Section IV-A, highlighting challenges in aligning with potentially noisy structured data
@@ IV-E Model-Expert Discrepancy Analysis
Our SHAP analysis (Figure 5) revealed that models often placed excessive weight on specific location-related terms without adequately considering the complete contextual narrative. We systematically quantified feature importance across all models, as shown in Table VI. The BSE model showed the highest reliance on direct intersection-related keywords, with terms like ”intersection” and ”signal” accounting for 42.6% of its total feature importance. In contrast, USE distributed feature importance more evenly, with the top five keywords accounting for only 27.3% of its total attribution, placing substantial weight on contextual modifiers like ”approaching” (13%), ”after” (6.1%), and ”before” (5.8%).
@@ IV-F1 LLM Evaluation Methodology
Each LLM was provided with identical prompts asking it to classify crash narratives as intersection-related or non-intersection-related. The prompt included general instructions about crash location classification, a brief explanation of intersection criteria, and the narrative text. To ensure fairness, we used the following standardized prompt structure:
@@ V Discussion and Conclusion
This study demonstrates an inverse relationship between deep learning model accuracy and expert judgment alignment in crash narrative classification. Our analysis of specialized DL models and LLMs reveals higher technical accuracy often correlates with lower expert agreement, while moderate-accuracy models better capture human reasoning patterns.
USE and BZS showed stronger expert alignment among specialized models, suggesting sentence-level embeddings better reflect holistic reasoning than BERT approaches that excel at pattern recognition but miss contextual nuance. LLMs achieved higher expert agreement ($\kappa$ up to 0.81) despite lower technical accuracy, with their high consensus (94.3%) indicating convergence on interpretations prioritizing context over keywords.
These discrepancies stem from fundamental differences: experts employ holistic reasoning considering temporal sequences and spatial relationships, while high-accuracy models rely on location-specific terms. SHAP analysis confirmed models with higher expert agreement distribute feature importance more evenly across contextual terms rather than fixating on keywords.
For transportation agencies, we recommend: (1) incorporate expert-alignment metrics alongside accuracy, (2) implement hybrid approaches leveraging complementary strengths, (3) establish confidence thresholds for human review, and (4) consider LLMs as cost-effective intermediaries for validation.
Future research should examine if this inverse accuracy-agreement relationship persists across different classification tasks, develop approaches combining high-accuracy models with expert-aligned LLMs, and investigate methods to distill contextual reasoning into efficient systems. As automated systems increasingly support transportation safety decisions, ensuring alignment between algorithmic classifications and expert judgment remains essential.